# Prompt-based Dynamic Token Pruning to Guide Transformer Attention in Efficient Segmentation

### v1

@@ 4 Conclusions and Discussion
This study introduces an adaptive framework for token pruning in Vision Transformers (ViT), designed to decrease the computational resources required to process irrelevant tokens. Using spatial priors through box prompts enhances the retention of semantically relevant tokens related to the target anatomical structure, thus improving segmentation accuracy. The calculation of the entropy-guided similarity score, to identify tokens aligned with the spatial prior, ensures that the token selection process is data-driven. This allows robustness to variations within the different input images. The experimental results suggest that the framework reduces the computational cost while maintaining (or even improving) the segmentation performance in certain scenarios. This study can be further extended to other transformer-based segmentation models and incorporate multiple types of prompts to gain additional insight into the generalizability of the proposed pruning approach.

### v2

@@ V Conclusion
This study introduces PrATo, an adaptive framework for token pruning in Vision Transformers (ViT), designed to decrease the computational resources required to process irrelevant tokens. Using spatial priors through box prompts enhances the retention of semantically relevant tokens related to the target anatomical structure, thus improving segmentation accuracy. The calculation of the entropy-guided similarity score, to identify tokens aligned with the spatial prior, ensures that the token selection process is data-driven. This allows robustness to variations within the different input images. The experimental results suggest that the framework offers highly accurate segmentation performance and faster inference speed than other approaches. This approach reduces computational cost while maintaining (or even improving) the segmentation performance of several ViT-based segmentation models. This demonstrates the applicability of the proposed framework across different architectures. PrATo shows robustness by demonstrating graceful degradation with inaccurate prompts. Thus, the proposed framework is highly relevant for automated medical image segmentation as it tackles the high computational overhead associated with ViTs. This makes them deployable in resource-constrained settings.
The performance of PrATo is highly dependent on the quality of the box-prompts. Furthermore, the dynamic hard pruning mechanism of PrATo may discard fine-grained features related to the boundaries of the anatomical structures. The proposed framework can be improved by integrating fine-grained prompts, such as points and scribbles, to enhance the boundary delineation capacity while preserving the efficient segmentation performance across a wide range of medical image analysis tasks and modalities.
VI Acknowledgments
This work was supported by the J. C. Bose National Fellowship, grant no. JCB/2020/000033 of S. Mitra. The authors acknowledge the use of computing resources provided by IDEAS-TIH, ISI Kolkata, India.